One number is 4 less than another. Their sum is 18. Find the numbers.
Answer by ikleyn(52805)

.

Let x be "another number".


Then "one number" is  (x-4),  and your equation is


    x + (x-4) = 18.


Simplify and solve


    2x = 18+4 = 22

     x        = 22/2 = 11.


ANSWER.  The numbers are 11 and (11-4) = 7.

Solved.